# PhishIntel: Toward Practical Deployment of Reference-based Phishing Detection

### v1

@@ Abstract.
Phishing is a critical cyber threat, exploiting deceptive tactics to compromise victims and cause significant financial losses. While reference-based phishing detectors (RBPDs) achieve high precision by analyzing brand-domain consistency, their real-world deployment is hindered by challenges such as high latency and inefficiency in URL analysis. To address these limitations, we present PhishIntel, an end-to-end phishing detection system for real-world deployment. PhishIntel intelligently determines whether a URL can be processed immediately or not, segmenting the detection process into two distinct tasks: a fast task that checks against local blacklists and result cache, and a slow task that conducts online blacklist verification, URL crawling, and webpage analysis using an RBPD. This fast-slow task system architecture ensures low response latency while retaining the robust detection capabilities of RBPDs for zero-day phishing threats. Furthermore, we develop two downstream applications based on PhishIntel: a phishing intelligence platform and a phishing email detection plugin for Microsoft Outlook, demonstrating its practical efficacy and utility.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Intrusion detection systemsccs: Software and its engineering Real-time systems software
@@ 2.1. Problem Statement
Phishing detection is essentially a binary classification task. Given a URL $u$, a phishing detector $\mathcal{D}$ analyzes the information associated with $u$, such as its webpage screenshot $s$ and HTML content $h$, and outputs a detection result $r$, classifying it as either benign or phishing.
In real-world scenarios, a deployed phishing detection system $\mathcal{F}$ typically processes a large volume of URLs. Formally, each incoming batch of URLs can be represented as a list $\mathcal{U}=\{u_{1},u_{2},...,u_{n}\}$, with the corresponding detection results denoted as =$\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$, where $n$ is the number of URLs in this batch. Their associated webpage screenshots $\mathcal{S}=\{s_{1},s_{2},...,s_{n}\}$ and HTML content =$\mathcal{H}=\{h_{1},h_{2},...,h_{n}\}$ are not initially available and must be fetched using an additional webpage crawler $\mathcal{C}$. However, integrating $\mathcal{C}$ introduces significant runtime overhead, leading to increased system latency. This latency is particularly problematic in real-world applications, such as phishing email detection, where high throughput and responsiveness are critical for a positive user experience.
@@ 2.4. Slow Task Worker
Specifically, an STW consists of three major components, an online blacklist checker, a webpage crawler, and an RBPD. The online blacklist checker initially checks $u$ against proprietary, non-publicly accessible blacklists via their URL matching APIs. This serves as an additional layer of the URL filtering process to complement FTW. If a match is found, the STW records a phishing result in the cache. Otherwise, the webpage crawler retrieves the webpage content of $u$, including its screenshot and HTML. This content is subsequently analyzed by the RBPD, which generates a detection result that is also stored in the result cache.
@@ 3.1. Phishing Intelligence Platform
This platform also features a dashboard page that visualizes phishing detection statistics derived from the result cache in PhishIntel, offering deeper insights into the current phishing attack landscape. For example, users can explore the number of unique phishing URLs, targeted entities, and recent trends, such as the daily detection count of phishing URLs and the most frequently targeted entities. Deployers can utilize this data to notify users about emerging phishing trends, safeguarding end users from such attacks through regular reminders and alerts.
@@ 3.2. Phishing Email Detection
We also integrate PhishIntel into an enterprise-level email system, as a phishing email detection plugin in Outlook. Specifically, we create a button in the tool menu. When users encounter a suspicious email and would like to report it for analysis, they can click this button to activate the phishing email detection tool. After clicking the Check button in the tool, it will automatically extract all the URLs in the selected email as a list, and send it to the PhishIntel API. For the particular case in Figure 2(b), the only URL in the email is classified as phishing, hence, the email is regarded as a phishing email.
@@ 5. Conclusion
In this paper, we introduce PhishIntel, an end-to-end phishing detection system designed for real-world deployment. PhishIntel integrates RBPDs, blacklists, and user reports collaboratively, enhancing both the effectiveness and efficiency of phishing detection. Its fast-slow task system architecture ensures real-time responsiveness and improves user experience. We demonstrate its practical utility through two applications: a phishing intelligence platform and a phishing email detection plugin in Microsoft Outlook. Looking ahead, we plan to optimize PhishIntel for larger-scale deployments, such as in telecommunication data centers.

### v2

@@ Abstract.
Phishing is a critical cyber threat, exploiting deceptive tactics to compromise victims and cause significant financial losses. While reference-based phishing detectors (RBPDs) have achieved notable advancements in detection accuracy, their real-world deployment is hindered by challenges such as high latency and inefficiency in URL analysis. To address these limitations, we present PhishIntel, an end-to-end phishing detection system for real-world deployment. PhishIntel intelligently determines whether a URL can be processed immediately or not, segmenting the detection process into two distinct tasks: a fast task that checks against local blacklists and result cache, and a slow task that conducts online blacklist verification, URL crawling, and webpage analysis using an RBPD. This fast-slow task system architecture ensures low response latency while retaining the robust detection capabilities of RBPDs for zero-day phishing threats. Furthermore, we develop two downstream applications based on PhishIntel: a phishing intelligence platform and a phishing email detection plugin for Microsoft Outlook, demonstrating its practical efficacy and utility.
@@
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715192isbn: 979-8-4007-1331-6/25/04ccs: Security and privacy Intrusion detection systemsccs: Software and its engineering Real-time systems software
@@ 2.1. Problem Statement
Phishing detection is essentially a binary classification task. Given a URL $u$, a phishing detector $\mathcal{D}$ analyzes the information associated with $u$, such as its webpage screenshot $s$ and HTML $h$, and outputs a detection result $r$, classifying it as either benign or phishing.
In real-world scenarios, a deployed phishing detection system $\mathcal{F}$ typically processes a large volume of URLs. Formally, each incoming batch of URLs can be represented as a list $\mathcal{U}=\{u_{1},u_{2},...,u_{n}\}$, with the corresponding detection results denoted as =$\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$, where $n$ is the number of URLs in this batch. Their associated webpage screenshots $\mathcal{S}=\{s_{1},s_{2},...,s_{n}\}$ and HTML =$\mathcal{H}=\{h_{1},h_{2},...,h_{n}\}$ are not initially available and must be fetched using an additional webpage crawler $\mathcal{C}$. However, integrating $\mathcal{C}$ introduces significant runtime overhead, leading to increased system latency. This latency is particularly problematic in real-world applications, such as phishing email detection, where high throughput and responsiveness are critical for a positive user experience.
@@ 2.4. Slow Task Worker
Specifically, an STW consists of three major components, an online blacklist checker, a webpage crawler, and an RBPD. The online blacklist checker initially checks $u$ against proprietary, non-publicly accessible blacklists via their URL matching APIs. This serves as an additional layer of the URL filtering process to complement FTW. If a match is found, the STW records a phishing result in the result cache. Otherwise, the webpage crawler retrieves the webpage content of $u$, including its screenshot and HTML. This content is subsequently analyzed by the RBPD, which generates a detection result that is also stored in the result cache.
@@ 3.1. Phishing Intelligence Platform
This platform also features a dashboard page that visualizes phishing detection statistics derived from the result cache in PhishIntel, offering deeper insights into the current phishing attack landscape. For example, users can explore the number of unique phishing URLs, targeted entities, and recent trends, such as the daily detection count of phishing URLs and the most frequently targeted entities. Deployers can utilize this data to notify users about emerging phishing trends, safeguarding them from such attacks through regular reminders and alerts.
@@ 3.2. Phishing Email Detection
We also integrate PhishIntel into an enterprise-level email system, as a phishing email detection plugin in Outlook. Specifically, we create a button in the tool menu. When users encounter a suspicious email and would like to report it for analysis, they can click this button to activate the phishing email detection tool. After clicking the Check button in the tool, it will automatically extract all the URLs in the selected email as a list, and send it to PhishIntel. For the particular case in Figure 2(b), the only URL in the email is classified as phishing, hence, the email is deemed as a phishing email.
@@ 5. Conclusion
In this paper, we introduce PhishIntel, an end-to-end phishing detection system designed for real-world deployment. PhishIntel integrates RBPDs, blacklists, and user reports collaboratively, enhancing both the effectiveness and efficiency of phishing detection. Its fast-slow task system architecture ensures real-time responsiveness and improves user experience. We demonstrate its practical utility through two applications: a phishing intelligence platform and a phishing email detection plugin in Microsoft Outlook. Looking ahead, we plan to optimize PhishIntel for larger-scale deployments.
Acknowledgements.
This work was supported in part by the National Research Foundation Singapore, NCS Pte. Ltd. and National University of Singapore under the NUS-NCS Joint Laboratory (Grant A-0008542-00-00).